OpenStudy (keana):

What are the x-intercept(s) of the graph of y - x^2 = -3x - 10? A:(2, 0) and (-5, 0) B:(-2, 0) and (5, 0) C:(-2, 0) and (-5, 0) D:(2, 0) and (5, 0)

OpenStudy (pfenn1):

Set y = 0 and solve for x.

OpenStudy (pfenn1):

\[0 - x^2 = -3x - 10?\]\[x^2-3x-10=0\]Either factor or use the quadratic formula to solve.

mathslover (mathslover):

we have to find x - intercept this means that y = 0 since this equation should lie on real axis hence putting y = 0 \[0-x^2=-3x-10\] \[-x^2=-3x-10\] \[x^2-3x-10=0\] \[x^2+2x-5x-10=0] \[x(x+2)-5(x+2)=0\] \[(x-5)(x+2)=0\] \[x=5 , x=-2\]

mathslover (mathslover):

hence it is (-2,0)(5,0)